# Preprint submitted to Mathematical Geosciences

### v1

@@ 5.2 History matching results for Case 1
Finally, we present prior and posterior distributions for the geological scenario parameters – mud fraction, channel orientation, and channel width. In Figure 15 we display the prior (uniform) distribution in gray and the posterior histograms in light blue. The vertical red lines indicate the true values. The prior ranges and true values are as given in Tables 1 and 2, respectively.
The CNN-based procedure described in Section 4 is used to infer the channel width and orientation for each posterior realization (facies fractions are computed directly), and the histograms are then constructed from the full set of posterior geomodels.
Very close alignment between the posterior parameter distributions and the true values is observed in Figure 15. Some amount of spread in the posterior histograms is evident, though the mode of each distribution is near the true scenario parameter in all cases. Consistent with the posterior geomodels discussed earlier, these results highlight the ability of the overall procedure to provide uncertainty reduction in the scenario parameters. We reiterate that this is achieved automatically within our workflow – we do not need to perform any sort of two-level history matching where the scenario parameters are determined in an initial step. This represents a significant advantage of our (unified) multiscenario parameterization.
@@
Figure 15: Case 1: history matching results for geological scenario parameters. Gray regions show the (uniform) prior distribution and blue histograms denote the posterior distribution. Vertical red dashed lines indicate true values.
@@ 5.3 Aggregate results for Cases 2 and 3
The prior and posterior scenario parameter distributions, along with the true values, appear in Figure 17. As was observed for Case 1 (Figure 15), we see substantial uncertainty reduction in these parameters for Cases 2 and 3. The posterior histograms include the true values in all cases, though slight deviation between the true values and the posterior modes is evident in Figure 17b,c,f. We do not view this as a concern since the modes of the marginal distributions need not coincide with the mode of the joint distribution in general multivariate problems.
@@ 6 Concluding remarks
In this study, we developed a latent diffusion model for the parameterization of large 3D systems with geological scenario uncertainty. Our 3D-LDM procedure provides a direct mapping from geomodel space to latent space (and vice versa), along with substantial dimension reduction. The implementation combines a 3D variational autoencoder and a 3D U-net, trained sequentially, and it includes a novel perceptual loss term that acts to improve the realism of the generated geomodels. An additional loss term enables conditioning to hard data. We applied the 3D-LDM for the parameterization of three-facies fluvial channel systems with variable (uncertain) scenario parameters. These included mud facies fraction, channel orientation, and channel width. For any set of scenario parameters, an essentially infinite number of realizations can be generated, so the range of uncertainty considered with the 3D-LDM procedure can be very wide.
@@
We then performed history matching for an immiscible, two-phase flow problem involving three injection wells and six production wells, all operating under BHP control. Uniform priors on the uncertain scenario parameters were applied. The history matching procedure entailed adjusting the latent variables, using ESMDA, to achieve low mismatch between simulation results and (synthetic) field observations. Three ‘true’ models characterized by different geological scenario parameters were tested. For all cases, we observed significant uncertainty reduction and a high degree of consistency between the posterior scenario parameters and the true parameters. Posterior geomodels were also shown to resemble the true model much more closely than the prior models. Predictions for well rates were presented for Case 1, and these showed uncertainty reduction and correspondence with the true results.
Future research directions include testing our parameterization method with more rigorous history matching algorithms, such as Markov chain Monte Carlo (MCMC), to assess the accuracy of the ESMDA procedure used here. The stochastic character of Monte Carlo-based methods may, in some respects, be more compatible with the 3D-LDM latent space, so MCMC can be used to provide benchmark results. Because MCMC typically requires ${\mathcal{O}}(10^{4}-10^{5})$ forward runs, a fast surrogate model for flow simulation will be needed for this assessment.
@@ Acknowledgments
We are grateful to the industrial affiliates of the Stanford Smart Fields Consortium for financial support. We also thank the SDSS Center for Computation for providing the computational resources used in this work.
Code availability statement
The source codes and datasets used in this work will be made available for download upon publication. The codes are based on the implementations provided in the https://github.com/huggingface/diffusers and https://github.com/Project-MONAI/GenerativeModels repositories. The Python libraries diffusers and monai are used in our implementation.
Appendix
7 3D-LDM model architecture
@@
Latent var. Sampling $\sim\mathcal{N}(\mathcal{E}_{\mu},\mathcal{E}_{\sigma}^{2})$ $\left(\frac{N_{x}}{8},\frac{N_{y}}{8},\frac{N_{z}}{8},1\right)$

### v2

@@ 5.2 History matching results for Case 1
Finally, we present prior and posterior distributions for the geological scenario parameters – mud fraction, channel orientation, and channel width. In Figure 15, we display the histograms for the prior and the posterior distributions in gray and light blue, respectively. The vertical red lines indicate the true values, as given in Table 2. Mud facies fractions (Figure 15a) are computed directly from the (prior or posterior) geomodel realizations. For channel width and orientation, the CNN-based procedure ($\tilde{\mathbf{s}}=\text{CNN}(\mathbf{r})$) described in Section 4 is applied for each prior and posterior realization. Histograms of the scenario parameters are then constructed from the full set of $N_{e}=200$ prior and posterior geomodels. Note that the histograms for the prior models are slightly nonuniform, in contrast to the specifications of Table 1. This is due to limitations and approximations in the OBM procedure applied to generate the training samples and to Monte Carlo error.
Very close alignment between the posterior parameter distributions and the true values is observed in Figure 15. Some amount of spread in the posterior histograms is evident, though the mode of each distribution is near the true scenario parameter in all cases. Consistent with the posterior geomodels discussed earlier, these results highlight the ability of the overall procedure to provide uncertainty reduction in the scenario parameters. We reiterate that this is achieved automatically within our workflow – we do not need to perform any sort of two-level history matching where the scenario parameters are determined in an initial step. This represents a significant advantage of our unified parameterization.
@@
Figure 15: Case 1: history matching results for geological scenario parameters. Gray regions show the prior distribution and blue histograms denote the posterior distribution. Vertical red dashed lines indicate true values.
@@ 5.3 Aggregate results for Cases 2 and 3
The prior and posterior scenario parameter distributions, along with the true values, appear in Figure 17. The priors again deviate from strict uniformity. Consistent with Case 1 (Figure 15), we see substantial uncertainty reduction in the scenario parameters for Cases 2 and 3. The posterior histograms include the true values in all cases, though slight deviation between the true values and the posterior modes is evident in Figure 17a,b,c. We do not view this as a concern since the modes of the marginal distributions need not coincide with the mode of the joint distribution in general multivariate problems.
@@ 6 Concluding remarks
In this study, we developed a latent diffusion model for the parameterization of large 3D systems with variable geological scenario parameters. Our 3D-LDM procedure provides a direct mapping from geomodel space to latent space (and vice versa), along with substantial dimension reduction. The implementation combines a 3D variational autoencoder and a 3D U-net, trained sequentially, and it includes a novel perceptual loss term that acts to improve the realism of the generated geomodels. An additional loss term enables conditioning to hard data. We applied the 3D-LDM for the parameterization of three-facies fluvial channel systems with variable (uncertain) scenario parameters. These included mud facies fraction, channel orientation, and channel width. For any set of scenario parameters, an essentially infinite number of realizations can be generated, so the range of uncertainty considered with the 3D-LDM procedure can be very wide.
@@
We then performed history matching for an immiscible, two-phase flow problem involving three injection wells and six production wells, all operating under BHP control. Approximately uniform priors on the uncertain scenario parameters were applied. The history matching procedure entailed adjusting the latent variables, using ESMDA, to achieve low mismatch between simulation results and (synthetic) field observations. Three ‘true’ models characterized by different geological scenario parameters were tested. For all cases, we observed significant uncertainty reduction and a high degree of consistency between the posterior scenario parameters and the true parameters. Posterior geomodels were also shown to resemble the true model much more closely than the prior models. Predictions for well rates were presented for Case 1, and these showed uncertainty reduction and correspondence with the true results.
Future research directions include testing our parameterization method with more rigorous history matching algorithms, such as Markov chain Monte Carlo (MCMC), to assess the accuracy of the ESMDA procedure used here. The stochastic character of Monte Carlo-based methods may, in some respects, be more compatible with the 3D-LDM latent space, so MCMC can be used to provide benchmark results. Because MCMC typically requires ${\mathcal{O}}(10^{4}-10^{5})$ forward runs, a fast surrogate model for flow simulation will be needed for this assessment. This surrogate will take the geomodel as input, and provide as output the injection and production rates. Research in this direction, using convolutional layers to extract spatial features and recurrent layers to capture temporal dynamics, is underway.
@@ Acknowledgments
We are grateful to the Stanford Doerr School of Sustainability and to the industrial affiliates of the Stanford Smart Fields Consortium for financial support. We also thank the SDSS Center for Computation for providing the computational resources used in this work.
Code availability
The source codes and datasets used in this work are available for download at https://github.com/guidodf09/ldm_3d_geomodel. The codes are based on the implementations provided in the https://github.com/huggingface/diffusers and https://github.com/Project-MONAI/GenerativeModels repositories. The Python libraries diffusers and monai are used in our implementation.
Appendix A
3D-LDM algorithms
Algorithm 1 VAE training
1:while not converged do
2:  Sample from dataset $\mathbf{m}_{0}\sim q(\mathbf{m}_{0})$
3:  Encode and decode sample $\bm{\xi}_{0}=\mathcal{E}(\mathbf{m}_{0}),\hat{\mathbf{m}}_{0}=\mathcal{D}(\bm{\xi}_{0})$
4:  Compute $L_{\text{VAE}}$ (Eq. 4), with:
5:   $L_{\text{recon}}=\|\mathbf{m}_{0}-\hat{\mathbf{m}}_{0}\|_{2}^{2}$
6:   $L_{\text{KL}}=D_{\text{KL}}\left[\mathcal{N}(\mathcal{E}_{\mu},\mathcal{E}_{\sigma}^{2})\,\|\,\mathcal{N}(\mathbf{0},\mathbf{I}_{n_{c}})\right]$
7:   $L_{\text{h}}=\frac{1}{N_{\text{h}}}\|\mathbf{H}(\mathbf{m}_{0}-\hat{\mathbf{m}}_{0})\|_{2}^{2}$
8:   $L_{\text{perc}}=\sum_{d\in\{x,y,z\}}\frac{1}{N_{d}^{\text{sub}}}\sum_{i=1}^{N_{d}^{\text{sub}}}\|\phi(\mathbf{m}_{0,i}^{(d)})-\phi(\hat{\mathbf{m}}_{0,i}^{(d)})\|_{2}^{2}$
9:  Gradient descent step on $\nabla_{\theta}L_{\text{VAE}}(\theta)$
10:end while
Algorithm 2 U-net training
1:while not converged do
2:  Sample from dataset $\mathbf{m}_{0}\sim q(\mathbf{m}_{0})$
3:  Encode sample $\bm{\xi}_{0}=\mathcal{E}(\mathbf{m}_{0})$
4:  Sample discrete step $t\sim U({1,\ldots,T})$ and noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n_{c}})$
5:  Compute $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$
6:  Perform noising step to compute $\bm{\xi}_{t}$ (Eq. 7)
7:  Predict $\bm{\epsilon}_{\theta}(\bm{\xi}_{t},t)$ with U-net
8:  Compute $L_{\text{U-net}}$ (Eq. 9)
9:  Gradient descent step on $\nabla_{\theta}L_{\text{U-net}}(\theta)$
10:end while
Algorithm 3 Inference (generation)
1:Sample noise $\bm{\xi}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n_{c}})$
2:for $t=T,\ldots,1$} do
3:  Compute $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$
4:  Predict $\bm{\epsilon}_{\theta}(\bm{\xi}_{t},t)$ with U-net
5:  Perform denoising step to compute $\bm{\xi}_{t-1}$ (Eq. 8)
6:end for
7:Decode denoised latent ${\mathbf{m}}_{0}^{\text{LDM}}=\mathcal{D}(\bm{\xi}_{0})$
8:return generated sample ${\mathbf{m}}_{0}^{\text{LDM}}$
Appendix B
3D-LDM model architecture
@@
Latent variable Sampling $\sim\mathcal{N}(\mathcal{E}_{\mu},\mathcal{E}_{\sigma}^{2})$ $\left(\frac{N_{x}}{8},\frac{N_{y}}{8},\frac{N_{z}}{8},1\right)$